# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ D.2 PG-DPO Implementation Details
  • Policy Network ($\bm{\pi}_{\theta}$): The policy network architecture consists of an input layer taking the current wealth $W_{t}$, current time $t$, and state vector $\mathbf{Y}_{t}$ (total $2+k$ inputs), followed by three hidden layers with 200 units each and LeakyReLU activation, and an output layer producing the $n$-dimensional portfolio weight vector $\bm{\pi}_{\theta}$.
@@ D.4 Deep BSDE Implementation Details
    • Y0Net: Approximates the value function at the initial time of a path, $Y_{t_{0}}\approx V(t_{0},X_{t_{0}},\mathbf{Y}_{t_{0}})$. It takes the initial state $(t_{0},\log X_{t_{0}},\mathbf{Y}_{t_{0}})$ (total $2+k$ inputs) as input and produces a scalar output (with a final negative softplus transformation). It typically uses 2 hidden layers of 64 units.
@@
  • Policy Derivation: The optimal portfolio $\bm{\pi}_{t}$ is dynamically constructed at each step. This involves using the network outputs $Y_{t}$ (obtained by evolving $Y_{t_{0}}$ via the BSDE discretization using ZNet) and $\bm{Z}_{t}$ (from ZNet) to approximate the first-order conditions of the HJB equation. This structure allows the policy to learn both myopic demand (from current factor states $\mathbf{Y}_{t}$) and intertemporal hedging demand (where terms involving $\bm{Z}^{Y}_{t}/Y_{t}$ approximate the necessary factor sensitivities, akin to the analytical solution’s form).
  • Loss Function: The Y0Net predicts $Y_{t_{0}}$. This value, along with ZNet’s outputs for $\bm{Z}_{t}$, is used to discretize the BSDE forward in time from $t_{0}$ to the terminal time $T$. The training objective is to minimize the mean squared error between the resulting $Y_{T}$ and the specified terminal utility $U(X_{T})=X_{T}^{1-\gamma}/(1-\gamma)$.

### v4

@@ D.2 PG-DPO Implementation Details
  • Policy Network ($\bm{\pi}_{\theta}$): The policy network consists of an input layer taking the current wealth $W_{t}$, current time $t$, and state vector $\mathbf{Y}_{t}$ (total $2+k$ inputs), followed by three hidden layers with 200 units each and LeakyReLU activation, and an output layer producing the $n$-dimensional portfolio weight vector $\bm{\pi}_{\theta}$.
@@ D.4 Deep BSDE Implementation Details
    • V0Net: Approximates the value function at the initial time of a path. Its output, denoted $\hat{V}_{t_{0}}$, is an approximation of the true value, $V(t_{0},X_{t_{0}},\mathbf{Y}_{t_{0}})$. It takes the initial state $(t_{0},\log X_{t_{0}},\mathbf{Y}_{t_{0}})$ as input and produces a scalar output (with a final negative softplus transformation). It typically uses 2 hidden layers of 64 units.
@@
  • Policy Derivation: The optimal portfolio $\bm{\pi}_{t}$ is not a direct network output but is dynamically constructed at each step. For CRRA utility, the FOC-derived policy simplifies significantly. The simplification stems from the value function’s known separable form in wealth (i.e., $V\propto X_{t}^{1-\gamma}$). This property allows the general intertemporal hedging demand term, $-\frac{1}{X_{t}V_{xx}}\Sigma^{-1}(\bm{\sigma\rho\sigma_{Y}})V_{x\mathbf{y}}$, to be analytically transformed into the more tractable expression $\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\bm{\rho}\right)\left(\bm{\sigma}_{Y}\frac{\nabla_{\mathbf{y}}V}{V}\right)$.
    The Deep BSDE implementation then creates a policy by combining an analytical myopic demand with a network-based approximation of this tractable hedging term. Specifically, the core theoretical component $\bm{\sigma}_{Y}\frac{\nabla_{\mathbf{y}}V}{V}$ is approximated by the ratio of neural network outputs, $\frac{\bm{Z}^{Y}_{t}}{\hat{V}_{t}}$. [^8] In our BSDE parameterization, the network ZNet learns $\bm{Z}^{Y}_{t}\approx\nabla_{\mathbf{y}}V\cdot\bm{\sigma}_{Y}$, where $\bm{\sigma}_{Y}$ is the diagonal matrix of factor volatilities. The ratio in the policy formula thus correctly approximates the required theoretical term $\frac{\nabla_{\mathbf{y}}V\cdot\bm{\sigma}_{Y}}{V}$. This leads to the final implemented formula:
    $\bm{\pi}_{t}=\underbrace{\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}\right)}_{\text{Myopic Demand}}\quad+\underbrace{\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\bm{\rho}\frac{\bm{Z}^{Y}_{t}}{\hat{V}_{t}}\right)}_{\text{Intertemporal Hedging Demand}}.$ (28)
    This structure is explicitly designed to capture both demand components, but as shown in our numerical results, the accuracy of this network-based approximation can be a practical point of failure.
  • Loss Function: The training objective is to accurately match the BSDE’s terminal condition. For each sample path starting at state $(t_{0},X_{t_{0}},\mathbf{Y}_{t_{0}})$, the ‘V0Net’ provides an initial value estimate, $\hat{V}_{t_{0}}$. This single initial estimate is then evolved forward to the terminal time $T$ using the discrete-time BSDE formula, which incorporates the outputs of the ‘ZNet’ at each intermediate time step. This process yields a simulated terminal value for that specific path, which we denote $\hat{V}_{T}^{\text{sim}}$. The loss is then calculated as the mean squared error between this simulated terminal value and the true terminal condition—the utility of the wealth $U(X_{T})$ realized at the end of that same path:
    $\mathcal{L}=\mathbb{E}\left[\left(\hat{V}_{T}^{\text{sim}}-U(X_{T})\right)^{2}\right].$
    The parameters of both ‘V0Net’ and ‘ZNet’ are optimized to minimize this terminal loss, effectively forcing the initial prediction $\hat{V}_{t_{0}}$ to be consistent with the terminal utility value through the lens of the BSDE dynamics.